### v1

@@ 9 MPD in self-supervised learning methods
Linear evaluation results of SimcLR [7] self-supervised MPD (ssl:MPD) and MPD IB (ssl:MPD IB) models, presented in Tables 12 and 13, align remarkably with the trends observed in their fine-tuned counterparts (Tables 6 and 7). This complementary relationship underscores the robustness of MPD models in both linear and fine-tuned scenarios. Specifically, ssl:MPD and ssl:MPD IB demonstrate impressive handling of perspectiveness in the ImageNet-PD subsets during linear evaluation, with ssl:MPD achieving an average Top-1 accuracy of 50.33±0.04, and ssl:MPD IB even more impressive at 51.62±0.03. These figures notably surpass the standard self-supervised model’s performance, which scores 35.08±0.04 and 32.53±0.04, respectively. This substantial improvement in handling perspective distortions is consistent with the fine-tuned results, where both ssl:MPD and ssl:MPD IB models maintained high accuracy levels, showing their efficacy in learning robust representations. Furthermore, the linear evaluation results reveal that these improvements are not at the expense of general performance. Both ssl:MPD and ssl:MPD IB sustain competitive accuracies on the standard ImageNet validation set, mirroring the fine-tuning outcomes where they demonstrated broad applicability and versatility. These findings from the linear evaluation and fine-tuning phases collectively reinforce the conclusion in our main paper about the enhanced capabilities of MPD models in self-supervised learning environments. They affirm the models’ ability to learn perspective-aware representations that are not only applicable to specialized scenarios involving perspective distortions but also effective across a wide range of visual recognition tasks. Linear evaluation results for ssl:MPD and ssl:MPD IB strongly reinforce the claim made in our main paper about their enhanced ability in self-supervised learning contexts. These models show not just incremental but significant improvements in handling complex visual distortions, establishing MPD as a powerful tool for self-supervised learning in real-world applications where perspective variations are prevalent.
@@
Table 12: Linear evaluation performance comparison for self-supervised models, standard SimCLR [7] and ssl:MPD model on original ImageNet validation set and ImageNet-PD subsets (without background interpolation). Self-supervised pretraining was performed on ImageNet dataset with batch size of 512 on linear learning rate for 100 epochs, refer SimCLR[7]. Top-1 accuracy reported. The probability of MPD is set to 0.8 for self-supervised pre-training and linear evaluation. ResNet50 is common backbone.
@@
Table 13: Linear evaluation performance comparison for self-supervised models, standard SimCLR [7] and ssl:MPD IB model on original ImageNet validation set and ImageNet-PD subsets (with background interpolation). Self-supervised pretraining was performed on batch size of 512 on linear learning rate for 100 epochs, refer SimCLR[7]. Top-1 accuracy reported. The probability of MPD IB is set to 0.8 for self-supervised pre-training and linear evaluation. ResNet50 is common backbone.
@@
MPD with DINO: Incorporating MPD into the knowledge distillation-based self-supervised method DINO [6] demonstrates MPD’s adaptability and effectiveness across different SSL approaches. The linear evaluation results for ssl:DINO-MPD and ssl:DINO-MPD IB, as detailed results presented in Tables 14 and 15, demonstrate the significant impact of integrating MPD and MPD IB into the DINO method. Both ssl:DINO-MPD and ssl:DINO-MPD IB models substantially improve handling perspectiveness on the ImageNet-PD subsets. Specifically, ssl:DINO-MPD achieves an average Top-1 accuracy of 60.70±0.04, while ssl:DINO-MPD IB further enhances this performance, achieving an even higher average Top-1 accuracy of 61.30±0.02. The integration of MPD and MPD IB into DINO demonstrates their adaptability and effectiveness in improving performance across self-supervised approaches, especially in scenarios involving perspectivness.
Table 14: Linear evaluation performance comparison for knowledge distillation based self-supervised method DINO [6]. Standard DINO is original method pretrained for 100 epochs with batch size of 512 and ssl: DINO-MPD model is MPD integrated DINO, following same pretraining hyperparameters. Post-pretraining, both the models are linearly trained for 100 epochs with batch size of 256. Top-1 accuracy reported. Evaluated on original ImageNet validation set and ImageNet-PD subsets (without background interpolation). The probability of MPD is set to 0.4 for self-supervised pre-training and 0.8 for linear evaluation. ViT-small transformer [10] is common backbone.
@@
Table 15: Linear evaluation performance comparison for DINO-MPD IB: Standard DINO [6] is original method pretrained for 100 epochs with batch size of 512 and ssl: DINO-MPD IB model is MPD IB integrated DINO, following same pretraining hyperparameters. Post-pretraining, both the models are linearly trained for 100 epochs with batch size of 256. Top-1 accuracy reported. Evaluated on original ImageNet validation set and ImageNet-PD subsets (with background interpolation). The probability of MPD IB is set to 0.4 for self-supervised pre-training and 0.8 for linear evaluation. ViT-small transformer [10] is common backbone.
@@
Figure 20: Activation maps of visual attention: Comparison of attention maps between DINO and MPD integrated DINO (ssl:DINO-MPD model). First attention head is used to visualize the attention. Example taken from ImageNet-PD set with original image. Both the models are pretrained for 100 epochs using ViT small architecture. DINO-MPD demonstrates on learning perspectiveness in input image.
Furthermore, the activation maps from ssl:DINO-MPD, as visualized in Fig. 20, qualitatively demonstrate its superior understanding of perspectiveness in input images. These maps reveal a more focused and relevant visual attention, indicating a deeper comprehension of the image structure and content, especially in perspectively distorted scenarios. The success of ssl:DINO-MPD in both quantitative performance and qualitative visual attention underscores the versatility of MPD. It shows that MPD can be effectively combined with various SSL methods, preserving their inherent strengths and augmenting them to handle better the challenges posed by perspective distortions. This enhances the claims made in our main paper about MPD’s capabilities, providing a strong case for its widespread applicability in diverse self-supervised learning approaches.
@@ 13.1 Crowd counting
Moreover, MPD-AutoCrowd further improves these metrics, achieving an MAE of 105.4 and an MSE of 124.61. These results underscore the capability of MPD to effectively deal with the complexities of crowd scenes, even when coupled with a different encoder. This ablation study complements the findings presented in Table 2 from the main paper, where MPD-CC and MPD-AutoCrowd outperform other crowd-counting methods across various benchmarks. The consistent improvement across different datasets and encoder configurations, especially with the challenging UCF-CC50 dataset, highlights MPD’s robustness and adaptability. MPD-AutoCrowd’s notable performance in handling varied crowd densities with high accuracy, particularly with an MAE of 96.80 and MSE of 139.50 on UCF-CC50, further validates the efficacy of MPD in specialized applications e.g., crowd counting. These results affirm the main paper’s assertion about the potential of MPD in complex visual recognition tasks and demonstrate its versatility in adapting to different encoder configurations while maintaining high performance.
@@ 13.2 Transfer learning on Fisheye images
Complementing these results, Tables 17 and 18 provide further insights into the models’ performances across a spectrum of MPD probabilities. The data reveal an optimal probability threshold for the MPD transformation application, which maximizes label efficiency. Notably, a probability setting of 0.5 for MPD yields the highest accuracy when all labels are utilized, indicating a non-linear relationship between the transformation probability and classification accuracy.
@@
The results substantiate the versatility and adaptability of MPD across different learning approaches and datasets, which is pivotal for real-world applications where perspective distortion is a significant challenge.
@@ 13.3 Person Re-Identification
Table 19: Ablation on the person re-identification for MPD (Clip-ReIdent) with CNN backbone: Explores multiple configurations of perceptiveness components with ResNet50x16 encoder. MPD probability set to 0.1, same as before. Min. (component) and Max. (component) defines the range of randomly incorporated perspectiveness. Component selection takes place between c-real and c-imaginary, with equal probability. Direction of each component also chosen with equal probability being positive or negative.
@@
14 Reproducibility and Further Research
We are committed to open research; we will release source code and models on GitHub and new benchmark dataset ImageNet-PD on public web-portal (link) to promote further.

### v2

@@ 9 MPD in self-supervised learning methods
Linear evaluation results of SimcLR [7] self-supervised MPD (ssl:MPD) and MPD IB (ssl:MPD IB) models, presented in Tables 13 and 14, align remarkably with the trends observed in their fine-tuned counterparts (Tables 7 and 8). This complementary relationship underscores the robustness of MPD models in both linear and fine-tuned scenarios. Specifically, ssl:MPD and ssl:MPD IB demonstrate impressive handling of perspective distortion in the ImageNet-PD subsets during linear evaluation, with ssl:MPD achieving an average Top-1 accuracy of 50.33±0.04, and ssl:MPD IB even more impressive at 51.62±0.03. These figures notably surpass the standard self-supervised model’s performance, which scores 35.08±0.04 and 32.53±0.04, respectively. This substantial improvement in handling perspective distortions is consistent with the fine-tuned results, where both ssl:MPD and ssl:MPD IB models maintained high accuracy levels, showing their efficacy in learning robust representations. Furthermore, the linear evaluation results reveal that these improvements are not at the expense of general performance. Both ssl:MPD and ssl:MPD IB sustain competitive accuracies on the standard ImageNet validation set, mirroring the fine-tuning outcomes where they demonstrated broad applicability and versatility. These findings from the linear evaluation and fine-tuning phases collectively reinforce the conclusion in our main paper about the enhanced capabilities of MPD models in self-supervised learning environments. They affirm the models’ ability to learn perspective-aware representations that are not only applicable to specialized scenarios involving perspective distortions but also effective across a wide range of visual recognition tasks. Linear evaluation results for ssl:MPD and ssl:MPD IB strongly reinforce the claim made in our main paper about their enhanced ability in self-supervised learning contexts. These models show not just incremental but significant improvements in handling complex visual distortions, establishing MPD as a powerful tool for self-supervised learning in real-world applications where perspective variations are prevalent.
@@
Table 13: Linear evaluation performance comparison for self-supervised models, standard SimCLR [7] and ssl:MPD model on original ImageNet validation set and ImageNet-PD subsets (black background). Self-supervised pretraining was performed on ImageNet dataset with batch size of 512 on linear learning rate for 100 epochs, refer SimCLR[7]. Top-1 accuracy reported. The probability of MPD is set to 0.8 for self-supervised pre-training and linear evaluation. ResNet50 is common backbone.
@@
Table 14: Linear evaluation performance comparison for self-supervised models, standard SimCLR [7] and ssl:MPD IB model on original ImageNet validation set and ImageNet-PD subsets (with integrated padding background). Self-supervised pretraining was performed on batch size of 512 on linear learning rate for 100 epochs, refer SimCLR[7]. Top-1 accuracy reported. The probability of MPD IB is set to 0.8 for self-supervised pre-training and linear evaluation. ResNet50 is common backbone.
@@
MPD with DINO: Incorporating MPD into the knowledge distillation-based self-supervised method DINO [6] demonstrates MPD’s adaptability and effectiveness across different SSL approaches. The linear evaluation results for ssl:DINO-MPD and ssl:DINO-MPD IB, as detailed results presented in Tables 15 and 16, demonstrate the significant impact of integrating MPD and MPD IB into the DINO method. Both ssl:DINO-MPD and ssl:DINO-MPD IB models substantially improves robustness against perspective distortion when evaluated on the ImageNet-PD subsets. Specifically, ssl:DINO-MPD achieves an average Top-1 accuracy of 60.70±0.04, while ssl:DINO-MPD IB further enhances this performance, achieving an even higher average Top-1 accuracy of 61.30±0.02. The integration of MPD and MPD IB into DINO demonstrates their adaptability and effectiveness in improving performance across self-supervised approaches, especially in scenarios involving perspective distortion.
Table 15: Linear evaluation performance comparison for knowledge distillation based self-supervised method DINO [6]. Standard DINO is original method pretrained for 100 epochs with batch size of 512 and ssl: DINO-MPD model is MPD integrated DINO, following same pretraining hyperparameters. Post-pretraining, both the models are linearly trained for 100 epochs with batch size of 256. Top-1 accuracy reported. Evaluated on original ImageNet validation set and ImageNet-PD subsets (black background). The probability of MPD is set to 0.4 for self-supervised pre-training and 0.8 for linear evaluation. ViT-small transformer [13] is common backbone.
@@
Table 16: Linear evaluation performance comparison for DINO-MPD IB: Standard DINO [6] is original method pretrained for 100 epochs with batch size of 512 and ssl: DINO-MPD IB model is MPD IB integrated DINO, following same pretraining hyperparameters. Post-pretraining, both the models are linearly trained for 100 epochs with batch size of 256. Top-1 accuracy reported. Evaluated on original ImageNet validation set and ImageNet-PD subsets (with integrated padding background). The probability of MPD IB is set to 0.4 for self-supervised pre-training and 0.8 for linear evaluation. ViT-small transformer [13] is common backbone.
@@
Figure 22: Activation maps of visual attention: Comparison of attention maps between DINO and MPD integrated DINO (ssl:DINO-MPD model). First attention head is used to visualize the attention. Example taken from ImageNet-PD set with original image. Both the models are pretrained for 100 epochs using ViT small architecture. DINO-MPD demonstrates on learning perspective distortion in input image.
Furthermore, the activation maps from ssl:DINO-MPD, as visualized in Fig. 22, qualitatively demonstrate its superior understanding of perspective distortion in input images. These maps reveal a more focused and relevant visual attention, indicating a deeper comprehension of the image structure and content, especially in perspectively distorted scenarios. The success of ssl:DINO-MPD in both quantitative performance and qualitative visual attention underscores the versatility of MPD. It shows that MPD can be effectively combined with various SSL methods, preserving their inherent strengths and augmenting them to handle better the challenges posed by perspective distortions. This enhances the claims made in our main paper about MPD’s capabilities, providing a strong case for its widespread applicability in diverse self-supervised learning approaches.
13 MPD vs. other augmentations
Table 17 demonstrates the effectiveness of MPD against other popular augmentation methods on the ImageNet validation set, ImageNet-PD subsets, and ImageNet-E [28]. Augmentation methods such as Mixup [58], Cutout [12], AugMix [17], and Pixmix [18] showed slight improvements on the ImageNet-PD benchmarks. For instance, Mixup improved the Top-1 accuracy to 65.46% on PD-T and 68.43% on PD-R, but these methods generally degraded performance on ImageNet-E, with Mixup dropping to 82.38%. This suggests that these augmentations do not adequately capture the characteristics of perspective distortions. In contrast, MPD-trained models, both in supervised and self-supervised settings, consistently outperformed all other methods on the perspective distortion benchmarks and ImageNet-E. The supervised MPD model achieved impressive Top-1 accuracies of 73.00%, 72.31%, 73.50%, and 72.91% on PD-T, PD-B, PD-L, and PD-R, respectively, and 85.92% on ImageNet-E. The self-supervised MPD model further enhanced performance, achieving Top-1 accuracies of 73.23%, 73.06%, 73.55%, and 73.54% on PD-T, PD-B, PD-L, and PD-R, respectively, and 86.66% on ImageNet-E. These results indicate that MPD effectively addresses perspective distortions, maintaining high accuracy across both the original and distorted datasets.
Table 17: Comparisons of MPD with other popular augmentation methods. ResNet50: baseline model trained on ImageNet [11] and subsequent models incorporates mentioned augmentation methods. All the models are evaluated on original ImageNet validation set, ImageNet-PD subsets (black background), and Imagenet-E. Both MPD trained models (supervised and SimCLR [7] self-supervised) outperformed on perspective distortion benchmark ImageNet-PD and ImageNet-E while maintain performance on original ImageNet validation set.
Method Original Val set ImageNet PD ImageNet-E
Top-view
(PD-T)
Bottom-view
(PD-B)
Left-view
(PD-L)
Right-view
(PD-R)
Top1 Top1/Top5 Top1/Top5 Top1/Top5 Top1/Top5 Top1
ResNet50 [40] 76.13 63.37/83.61 61.15/81.86 65.20/85.13 65.84/85.64 84.28
+Mixup [58] 77.46 65.46/85.15 66.79/85.91 68.02/86.98 68.43/87.22 82.38
+Cutout [12] 77.08 64.27/84.20 62.04/82.86 65.45/85.40 65.61/85.46 79.58
+AugMix [17] 77.53 64.12/84.01 62.90/82.33 65.95/85.94 66.49/86.03 80.97
+Pixmix [18] 77.37 65.52/85.21 64.76/84.54 67.26/86.67 67.56/86.80 82.38
+MPD 76.34 73.00/90.69 72.31/90.81 73.50/91.33 72.91/91.29 85.92
+MPD
76.77 73.23/91.39 73.06/91.54 73.55/91.53 73.54/91.65 86.66
@@ 14.1 Crowd counting
Moreover, MPD-AutoCrowd further improves these metrics, achieving an MAE of 105.4 and an MSE of 124.61. These results underscore the capability of MPD to effectively deal with the complexities of crowd scenes, even when coupled with a different encoder. This ablation study complements the findings presented in Table 2 from the main paper, where MPD-CC and MPD-AutoCrowd outperform other crowd-counting methods across various benchmarks. The consistent improvement across different datasets and encoder configurations, especially with the challenging UCF-CC50 dataset, highlights MPD’s robustness and adaptability. MPD-AutoCrowd’s notable performance in handling varied crowd densities with high accuracy, particularly with an MAE of 96.80 and MSE of 139.50 on UCF-CC50, further validates the efficacy of MPD in specialized applications, e.g., crowd counting.
Further, We adapted three augmentation methods tailored for crowd counting and observed MPD outperforms them, as shown in Table 19. Specifically, we compared MPD with RandAugment [10], OpenCV PT [59], and AugLy PT [38] across three datasets: SHHA, SHHB, and UCF_CC_50. MPD consistently achieved lower MAE and MSE values, indicating superior performance in handling crowd counting tasks. For instance, MPD-CC achieved a MAE of 51.93 and MSE of 84.3 on SHHA, significantly outperforming AugLy PT’s 56.18 MAE and 107.67 MSE. Furthermore, the samples shown in Fig. 23 illustrates that perspective transform augmentations from [38] and [59] do not capture the non-linearity inherent in real-world perspective distortions, as they primarily rely on affine transformations or translations. This limitation underscores the effectiveness of MPD in capturing complex distortions, leading to its superior performance in crowd counting benchmarks.
Table 19: Comparisons on augmentation methods in crowd counting. PT: Perspective Transform. Metrics: MAE and MSE errors
Method SHHA SHHB UCF_CC_50
MAE MSE MAE MSE MAE MSE
RandAugment [10] 59.81 113.48 10.03 19.84 187.62 271.28
OpenCV PT [59] 57.73 109.82 7.78 13.41 181.05 263.52
AugLy PT [38] 56.18 107.67 8.09 14.05 174.74 259.08
MPD-CC 51.93 84.3 6.73 9.82 101.3 140.65
MPD-AutoCrowd 50.81 85.01 6.61 9.58 96.8 139.5
Figure 23: Applied augmentations on crowd counting
These results affirm the main paper’s assertion about the potential of MPD in complex visual recognition tasks and demonstrate its versatility in adapting to different encoder configurations while maintaining high performance.
@@ 14.2 Transfer learning on Fisheye images
Complementing these results, Tables 20 and 21 provide further insights into the models’ performances across a spectrum of MPD probabilities. The data reveal an optimal probability threshold for the MPD transformation application, which maximizes label efficiency. Notably, a probability setting of 0.5 for MPD yields the highest accuracy when all labels are utilized, indicating a non-linear relationship between the transformation probability and classification accuracy.
@@
The results substantiate MPD’s versatility and adaptability across different learning approaches and datasets, which is pivotal for real-world applications where perspective distortion is a significant challenge.
@@ 14.3 Person Re-Identification
Table 22: Ablation on the person re-identification for MPD (Clip-ReIdent) with CNN backbone: Explores multiple configurations of perceptiveness components with ResNet50x16 encoder. MPD probability set to 0.1, same as before. Min. (component) and Max. (component) defines the range of randomly incorporated perspective distortion. Component selection takes place between c-real and c-imaginary, with equal probability. Direction of each component also chosen with equal probability being positive or negative.
@@
14.4 Object Detection
Table 23 presents the object detection performance of MPD-OD on the COCO [30] dataset, with varying probabilities of applying MPD-OD during training. Figure 25 shows more MPD-OD transformed image examples. The results show that as the probability of applying MPD-OD increases from 0.1 to 0.5, there is a consistent improvement in detection performance for both FasterRCNN [43] and FCOS [47] models. Specifically, FasterRCNN gradually increases IoU=0.50:0.95 from 37.60 to 40.00 and IoU=0.50 from 58.10 to 61.10. Similarly, FCOS shows an improvement from 39.10 to 40.20 in IoU=0.50:0.95 and from 58.50 to 60.30 in IoU=0.50. These trends indicate that higher probabilities of applying MPD-OD during training lead to better object detection performance, highlighting the effectiveness of MPD-OD in enhancing model robustness and accuracy. The consistent performance gains across different probabilities demonstrate the adaptability and efficacy of MPD-OD in handling diverse training scenarios and further strengthen the general usefulness of MPD.
Table 23: Object detection performance of MPD-OD on COCO [30] dataset. Ablations on probability to apply MPD-OD on examples during training. Metrics: IoU=0.50:0.95 / IoU=0.50
Method Original MPD-OD
P=0.1 P=0.2 P=0.3 P=0.4 P=0.5
FasterRCNN[43] (NeurIPS’15) 37.10/55.80 37.60/58.10 39.20/59.40 39.50/60.60 39.70/60.80 40.00/61.10
FCOS[47] (ICCV’19) 38.60/57.40 39.10/58.50 39.50/59.10 39.90/59.70 40.00/59.90 40.20/60.30
Figure 25: MPD transformed input & bounding boxes during training